The owner of a small hotel with 5 rooms is considering buying TV sets to rent to room occupants.

He expects that about half of his customers would be willing to rent sets and finally he buys 3 sets.

Assuming 100% occupancy at all times

i) What fraction of evenings will there be more requests than TV sets?

ii) What is the probability that a customer who requests a TV set will receive one?


1
Expert's answer
2021-10-25T18:56:15-0400

Let "X=" the number of requests: "X\sim Bin (n, p)"

Given "n=5, p=0.5, q=1-p=0.5"

i)


"P(X>3)=P(X=4)+P(X=5)"

"=\dbinom{5}{4}(0.5)^4(0.5)^{5-4}+\dbinom{5}{5}(0.5)^5(0.5)^{5-5}=0.1875"


ii)


"P(X\leq3)=1-P(X=4)-P(X=5)"

"=1-\dbinom{5}{4}(0.5)^4(0.5)^{5-4}-\dbinom{5}{5}(0.5)^5(0.5)^{5-5}"

"=0.8125"